Uranus beta is 1.2 and market risk premium 7 percent risk free rate is 10 percent. Uranus has debt is 14 percent tax rate is 35% what is WACc
1
Expert's answer
2021-07-29T10:55:01-0400

The capital asset pricing model is used for calculating the value of the cost of equity. The formula for the calculation cost of equity using CAPM is:

"Re = Rf + \u03b2 \u00d7 (Rm \u2212 Rf)"

Where:

Rf = the risk-free rate 

β = equity beta 

Rm = annual return of the market


The second step is calculating the cots of debt and preference share capital. It is very easy and simple to calculate the cost of debt. We need just need to know the cost and tax rate. Interest payments are tax-deductible. The cost of debt needs to be multiplied by (1 – tax rate), for the calculations for the weighted average cost of capital.

"Re = Rf + \u03b2 \u00d7 (Rm \u2212 Rf)"

Where:  

Re= cost of equity

Rf = the risk−free rate  

β = equity beta  

Rm = annual return of the market


"Re = Rf + β × (Rm − Rf)\\

Re=10+1.2×7\\

Re=18.4"

Cost of debt = cost of debt(1−tax rate)

cost of debt "=14(1\u22120.35)"

cost of debt "=9.1"

Weighted average cost of capital= cost of equity + cost of debt

Weighted average cost of capital "=18.4+9.1=27.5"